# Scaf-GRPO: Scaffolded Group Relative Policy Optimization for Enhancing LLM Reasoning

### v1

@@ 4.4  Further Analysis
Scaf-GRPO succeeds by fostering skill acquisition, not just imitation. Figure 5 illustrates this on a true-hard problem. The model’s reasoning evolves from imitating concrete “Solution Hints” to following strategic “Planning Hints” and finally to applying abstract “Knowledge Hints.” The key evidence for true learning appears when the model later revisits the problem. Without any hints, it successfully solves it by combining skills from its prior guided experiences. This demonstrates that hints serve as a tool to build lasting abilities, not just to achieve temporary success. This process of skill-building on hard problems overcomes the learning cliff.
@@
To validate our data filtering strategy, we train models on both the complete dataset and our filtered subset. As detailed in Table 3, the harder, filtered data yields a marginal 0.5% relative gain for vanilla GRPO but a substantial 6.0% boost for Scaf-GRPO on Qwen2.5-Math-7B. This disparity underscores that exposing a model to difficult problems is insufficient. A challenging curriculum is most effective when paired with a robust learning framework like Scaf-GRPO, which converts these challenges into learning opportunities.
@@
To verify that Scaf-GRPO cultivates robust reasoning skills rather than in-domain pattern matching, we evaluate its generalization on the out-of-distribution GPQA-Diamond benchmark, which features expert-level scientific questions. Scaf-GRPO demonstrates strong OOD performance, elevating the Qwen2.5-Math-7B model’s pass@1 score to 37.3%. This represents a significant improvement over the base model (24.7%) and the vanilla GRPO baseline (32.3%), and matches the performance of the strong prefix-based LUFFY baseline. This result indicates that the problem-solving abilities fostered by our scaffolding approach are fundamental and transfer effectively to novel domains, a key attribute for building more general AI reasoners.
Table 4: OOD performance (pass@1,%) of Qwen2.5-Math-7B on the GPQA-Diamond benchmark.
@@
SimpleRL-Zero simplerlzoo 33.3
Oat-Zero liu2025there 33.3
LUFFY yan2025learning 37.3
@@ Conclusion.
Appendix A Dataset and Benchmark Details
A.1  Training Data Source and Filtering
Our training data is derived from the DeepScaleR-Preview-Dataset [21], a comprehensive collection of 40k mathematical problems. Its contents are sourced from AIME, AMC, MATH [26], Still [6], and Omni-MATH [31]. To maximize training efficiency and target problems most conducive to learning, we implement a dynamic data filtering strategy tailored to each model’s capabilities. This strategy categorizes problems based on the model’s initial performance, assessed through 8 independent sampling attempts for each problem. These samples are generated using nucleus sampling with a temperature of 1.0 (top-p=1.0, top-k=-1) and a maximum length of 2048 tokens for both the prompt and the response. For our LongCoT model, this response token was increased to 8192 tokens to accommodate its generation style. Based on the outcomes, problems are categorized as follows:
@@ A.1  Training Data Source and Filtering
This filtering process results in a final training dataset where approximately 50% of the problems are from the “Potentially Solvable” category and the remaining 50% are "Too Hard". This curated dataset is challenging yet tractable, maximizing the efficiency of the training process for both the baseline GRPO and our Scaf-GRPO framework.
A.2  Hint Generation for Training Data
The hierarchical hints ($H_{\text{knowledge}},H_{\text{planning}},H_{\text{solution}}$) are the cornerstone of Scaf-GRPO’s guidance mechanism. To create them, we perform a one-time, offline preprocessing step using the powerful DeepSeek-R1 model [3]. For each problem in our curated training set, we provide the model with both the problem statement and its ground-truth solution trace.
We then employ a highly structured prompt, detailed in Appendix D, which is engineered not only to decompose the solution into our three-tiered hierarchy but also to enforce a crucial internal structure. Specifically, the prompt compels the model to generate exactly four numbered, progressive items for each category. These items are designed to build upon one another, creating four distinct levels of guidance. For instance, the four items for an $H_{\text{planning}}$ hint might represent: (1) the first step of the plan, (2) the second step, (3) the third step, and (4) the fourth step. This structured, multi-level design within each hint category enables the fine-grained, progressive exploration central to our method. The entire process ensures a consistent and high-quality set of structured hints, and the scripts will be included in our code release.
A.3  Evaluation Benchmarks
@@
A.4  Motivation for Dataset Selection
Our choice of the DeepScaleR-Preview-Dataset for training is deliberate. This decision is motivated by two key factors. First, the dataset’s broad scope, encompassing problems of varying difficulty, provides the rich and diverse material necessary for our filtering strategy to be effective. Second, its successful application in prior research [6] establishes it as a robust and relevant foundation for training advanced reasoning models.
For evaluation, our selection of benchmarks is designed for a rigorous and multifaceted assessment. The suite primarily consists of challenging competition-level benchmarks (AIME24/25, AMC, MATH, OlympiadBench, Minerva) and a standardized national exam (GaoKao2023), covering a broad spectrum of mathematical reasoning. Crucially, to measure out-of-distribution (OOD) generalization, we include the GPQA-Diamond benchmark [28]. As GPQA consists of graduate-level questions whose style and domain are distinct from our training data, strong performance on this benchmark indicates that Scaf-GRPO fosters genuine reasoning skills rather than mere pattern memorization of the training distribution.
@@ A.4  Motivation for Dataset Selection
AIME24 American Invitational Mathematics Examination 2024 [23] 30
AIME25 American Invitational Mathematics Examination 2025 [24] 30
AMC American Mathematics Competitions 2023 [25] 25
MATH-500 A subset from the MATH test set [26] 500
GaoKao2023en Chinese National College Entrance Exam 2023 [22] 385
OlympiadBench Math Olympiad-level problems [27] 675
Minerva A specialized dataset to evaluate quantitative and scientific reasoning abilities [32] 272
GPQA-Diamond Expert-level questions across biology, physics, and chemistry [28] 198
@@
Appendix B Experimental Setup Details
B.1  Computing Infrastructure
  • Framework: Our implementation is built upon the verl (0.4.1) framework [29], a robust and efficient library designed for large-scale reinforcement learning with LLMs.
B.2  Hyperparameter Details
@@
B.3  Implementation of Baseline Methods
@@
To compare against the dominant prefix-continuation paradigm, we train LUFFY [2] using its official public implementation and its original, recommended hyperparameter settings. To ensure a fair comparison of methodological effectiveness, we train it on our high-quality filtered dataset. This directly contrasts their off-policy guidance with our in-prompt scaffolding on identical data.
@@
For other leading methods like Simple-RL [7] and Oat-Zero [20], we do not perform any implementation or retraining. Instead, we evaluate their officially released, publicly available model checkpoints directly. All reported results for these models are obtained by running them through our unified evaluation pipeline, ensuring a consistent and fair comparison against established state-of-the-art work.
B.4  Evaluation Metrics Details
@@
  • For all mathematical reasoning benchmarks, we employ the “symeval” library [33], specifically its EvaluatorMathBatch module, to determine correctness. This approach moves beyond simple string comparison by using a sophisticated pipeline that combines regular expressions for robust answer extraction with SymPy for symbolic mathematical evaluation. This allows for the accurate verification of complex answers, including matrices, intervals, and symbolic expressions.
  • For the out-of-distribution GPQA-Diamond benchmark, we utilize the EleutherAI’s lm-evaluation-harness [34] to ensure a fair and standardized assessment. This widely adopted framework provides a consistent testing environment for generative models. We use its implementation of the gpqa-diamond task to compute the pass@1 score, thereby maintaining metric consistency while leveraging a community-standardized evaluation harness for OOD generalization.
Appendix C Methodology Details
@@
C.1  The Progressive Exploration and Replacement Algorithm
@@
The algorithm leverages the pre-generated, four-level progressive hint structure, detailed in Appendix A.2. It systematically searches through the hint categories in order of decreasing abstraction ($H_{\text{knowledge}}\rightarrow H_{\text{planning}}\rightarrow H_{\text{solution}}$) and, within each category, through the four levels of increasing detail.
@@
Appendix D Prompt Design in Scaf-GRPO
@@
D.1  Hint Generation Prompt
@@
Prompt for hint injection **[ROLE & GOAL]** {internallinenumbers*} You are an expert AI assistant specializing in problem-solving methodology and knowledge engineering. Your task is to analyze a given problem and its ground-truth solution, and then generate a structured breakdown of the reasoning process with a high degree of granularity. **[INPUT]** I will provide you with a "Problem" and its "Ground-Truth Solution". **[INSTRUCTIONS]** {internallinenumbers*} Based on the provided input, you must generate exactly THREE components. For each component, you MUST generate **a minimum of 4 numbered items**. {internallinenumbers*} - If a natural breakdown results in fewer than 4 items, you must **subdivide the existing steps into more detailed, finer-grained sub-steps** to meet the requirement. For example, a single calculation step can be broken down into ’substituting values’, ’performing the operation’, and ’stating the result’. {internallinenumbers*} 1. **Planning Skeleton**: Extract a high-level planning skeleton. This should be a concise, ordered list of the key reasoning steps and the overall strategy used to reach the solution. Do not include detailed calculations, just the logical flow. Break it down into at least 4 detailed steps. {internallinenumbers*} 2. **Knowledge Components**: Identify at least 4 essential knowledge components (like facts, definitions, theorems, lemmas, or formulas) required to solve the problem. List each component clearly in a numbered list. {internallinenumbers*} 3. **Solution Breakdown**: Divide the original Ground-Truth Solution into a numbered list of semantically coherent steps or chunks.There should be at least 4 steps or chunks. Each item in the list should be a direct quote or a faithful summary of a part of the original solution text. **[OUTPUT FORMAT]** {internallinenumbers*} You MUST provide your response in the following structured format. Ensure each section contains at least 4 items. <PLANNING_SKELETON> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </PLANNING_SKELETON> <KNOWLEDGE_COMPONENTS> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </KNOWLEDGE_COMPONENTS> <SOLUTION_BREAKDOWN> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </SOLUTION_BREAKDOWN> **[EXAMPLE]** --- BEGIN EXAMPLE --- # ... (Example Problem and Solution are the same) --- END EXAMPLE --- **[EXPECTED OUTPUT FOR THE EXAMPLE]** (This example now demonstrates the required granularity with >= 4 items per section) <PLANNING_SKELETON> 1. Identify the geometric shape (right-angled triangle) and the goal (find the hypotenuse). 2. Recall the relevant mathematical theorem that connects the sides of a right-angled triangle. 3. Formulate the equation by substituting the given side lengths (base and height) into the theorem. 4. Execute the arithmetic calculation to find the square of the hypotenuse. 5. Perform the final step of taking the square root to isolate the length of the hypotenuse. </PLANNING_SKELETON> <KNOWLEDGE_COMPONENTS> 1. Theorem: Pythagorean Theorem (a² + b² = c² for a right-angled triangle). 2. Definition: Hypotenuse (The longest side of a right-angled triangle, opposite the right angle). 3. Concept: Right-angled Triangle (A triangle with one angle measuring 90 degrees). 4. Mathematical Operation: Square Root (The inverse operation of squaring a number). </KNOWLEDGE_COMPONENTS> <SOLUTION_BREAKDOWN> {internallinenumbers*} 1. To find the hypotenuse of a right-angled triangle, we can use the Pythagorean theorem, which states that a² + b² = c², where a and b are the lengths of the two shorter sides (legs) and c is the length of the hypotenuse. 2. The given values are a = 4 cm and b = 3 cm. 3. Substituting these into the formula gives: 4² + 3² = 16 + 9 = 25. 4. This means c² = 25. Taking the square root of both sides results in c = 5 cm. </SOLUTION_BREAKDOWN> --- **[TASK]** Now, please process the following problem and solution, strictly following all instructions. **Problem**: {question} **Ground-Truth Solution**: {solution}
D.2  Inject Hints Prompt

### v2

@@ 4.4 Further Analysis
Scaf-GRPO succeeds by fostering skill acquisition rather than simple imitation. Figure 5 illustrates this trajectory on a challenging problem: the model evolves from utilizing concrete “Solution Hints” to abstract “Knowledge Hints,” and finally to solving the problem without any support. To quantify this transition from dependency to autonomy, we track “skill graduation” events over the first 300 training steps. We define this metric to capture the learning breakthrough specific to each method: for Scaf-GRPO, we count problems transitioning from hint-dependent to autonomous success, while for the Vanilla baseline, we track the shift from total failure to success. As shown in Table 4, Scaf-GRPO consistently yields a significantly higher volume of graduations (e.g., +137.8% on Qwen2.5-Math-1.5B). This confirms that our method effectively converts temporary guidance into lasting, independent reasoning capabilities. By enabling such robust skill-building on hard problems, Scaf-GRPO effectively overcomes the learning cliff.
Table 4: Comparison of total “graduations” across different model backbones.
Model Vanilla Scaf-GRPO (Ours) Relative Increase
Qwen2.5-Math-1.5B 1,123 2,670 +137.8%
Qwen2.5-Math-7B 434 483 +11.3%
Qwen2.5-7B 464 723 +55.8%
DeepSeek-R1-Distill-1.5B 453 694 +53.2%
Llama-3.2-3B-Instruct 577 986 +70.9%
@@
To validate our data filtering strategy, we train models on both the complete dataset and our filtered subset. As detailed in Table 3, the harder, filtered data yields a marginal 0.5% relative gain for vanilla GRPO but a substantial 6.0% boost for Scaf-GRPO on Qwen2.5-Math-7B. This disparity underscores that exposing a model to difficult problems is insufficient. A challenging curriculum is most effective when paired with a robust learning framework like Scaf-GRPO, which converts these challenges into learning opportunities. We also observe performance is sensitive to the density of solvable problems. Specifically, discarding “Too Easy” samples and subsampling the “Potentially Solvable” category at 50% achieves the optimal balance, yielding a 10.2% relative gain over the full dataset (Table 12). This empirical evidence justifies our training data selection. All experimental details are provided in Appendix G.2.
@@
To verify that Scaf-GRPO cultivates robust reasoning skills beyond in-domain pattern matching, we evaluate its generalization on the expert-level OOD benchmark, GPQA-Diamond. As shown in Table 6, Scaf-GRPO significantly enhances the Qwen2.5-Math-7B model, achieving a 15.5% relative improvement over Vanilla GRPO and matching the strong LUFFY baseline. Furthermore, this gain effectively transfers to the general-purpose Qwen2.5-7B-Base, where our method yields a 7.5% relative increase over Vanilla GRPO and outperforms LUFFY. These results demonstrate that the problem-solving abilities fostered by Scaf-GRPO are fundamental and agnostic to the model backbone.
Computational efficiency.
Scaf-GRPO optimizes training efficiency by applying guidance selectively and accelerating convergence. Empirically, during the training of Qwen2.5-Math-7B, the hint-guided exploration is triggered for only 17.4% of the samples, ensuring that the majority of the computational throughput remains dedicated to standard generation. More importantly, by converting zero-reward signals into high-value learning signals, Scaf-GRPO significantly reduces the total training duration required to reach optimal performance. As shown in Table 5, Scaf-GRPO reaches its best-performing checkpoint (50.9% avg.) in approximately 12 hours. This represents a clear efficiency gain over the Vanilla GRPO baseline, which requires 13 hours to reach a lower peak performance (45.2%), demonstrating that Scaf-GRPO achieves better results with a smaller total time budget.
Table 5: Training efficiency comparison on Qwen2.5-Math-7B. Best results are in bold.
Method Best Avg. (%) Time to Best Ckpt. Trigger Rate Peak Memory
Qwen2.5-Math-7B
Vanilla GRPO 45.2 $\sim$13 hours N/A $\sim$72 GB
Scaf-GRPO (Ours) 50.9 $\sim$12 hours 17.4% $\sim$73 GB
Table 6: OOD performance (pass@1,%) of Qwen2.5-Math-7B and Qwen2.5-7B-Base on the GPQA-Diamond benchmark.
@@
SimpleRL-Zero Zeng et al. (2025) 33.3
Oat-Zero Liu et al. (2025a) 33.3
LUFFY Yan et al. (2025) 37.3
@@
Qwen2.5-7B-Base
Vanilla GRPO 33.3
LUFFY Yan et al. (2025) 34.4
Scaf-GRPO (Ours) 35.8
@@ Conclusion.
Ethics Statement
This research adheres to the principles outlined in the ICLR Code of Ethics. Our primary objective is to contribute to society by advancing the frontiers of machine reasoning, upholding the highest standards of scientific excellence through transparent and reproducible methods. We acknowledge the potential for dual-use applications and the risks associated with the misuse of advanced AI reasoning systems. We also recognize the environmental impact of training large models and have striven for computational efficiency. Our work is based on publicly available datasets devoid of personal identifiable information. We are committed to fostering responsible innovation and encourage continued investigation into the societal impacts of increasingly capable models.
Reproducibility Statement
The supplementary material contains the complete source code to ensure full reproducibility. This includes the implementation of Scaf-GRPO and all scripts for data filtering and the training pipeline.
Acknowledgements
This work was supported in part by the Research Grants Council under the Areas of Excellence scheme grant AoE/E-601/22-R.
Appendix A The Use of Large Language Models (LLMs)
During the preparation of this manuscript, the authors employed a large language model (LLM) to aid in refining the language and correcting grammatical errors. The role of the LLM was strictly limited to that of a writing-enhancement tool. The authors take full responsibility for the substantive content, arguments, and final phrasing presented in this paper.
Appendix B Dataset and Benchmark Details
B.1 Training Data Source and Filtering
Our training data is derived from the DeepScaleR-Preview-Dataset (Luo et al., 2025), a comprehensive collection of 40k mathematical problems. Its contents are sourced from AIME, AMC, MATH (Hendrycks et al., 2021), Still (Min et al., 2024), and Omni-MATH (Gao et al., 2024a). To maximize training efficiency and target problems most conducive to learning, we implement a dynamic data filtering strategy tailored to each model’s capabilities. This strategy categorizes problems based on the model’s initial performance, assessed through 8 independent sampling attempts for each problem. These samples are generated using nucleus sampling with a temperature of 1.0 (top-p=1.0, top-k=-1) and a maximum length of 2048 tokens for both the prompt and the response. For our LongCoT model, this response token was increased to 8192 tokens to accommodate its generation style. Based on the outcomes, problems are categorized as follows:
@@ B.1 Training Data Source and Filtering
This filtering process results in a final training dataset where approximately 50% of the problems are from the “Potentially Solvable” category and the remaining 50% are ”Too Hard”. This curated dataset is challenging yet tractable, maximizing the efficiency of the training process for both the baseline GRPO and our Scaf-GRPO framework.
B.2 Hint Generation for Training Data
The hierarchical hints ($H_{\text{knowledge}},H_{\text{planning}},H_{\text{solution}}$) are the cornerstone of Scaf-GRPO’s guidance mechanism. To create them, we perform a one-time, offline preprocessing step using the powerful DeepSeek-R1 model (Guo et al., 2025). For each problem in our curated training set, we provide the model with both the problem statement and its ground-truth solution trace.
We then employ a highly structured prompt, detailed in Appendix E, which is engineered not only to decompose the solution into our three-tiered hierarchy but also to enforce a crucial internal structure. Specifically, the prompt compels the model to generate exactly four numbered, progressive items for each category. These items are designed to build upon one another, creating four distinct levels of guidance. For instance, the four items for an $H_{\text{planning}}$ hint might represent: (1) the first step of the plan, (2) the second step, (3) the third step, and (4) the fourth step. This structured, multi-level design within each hint category enables the fine-grained, progressive exploration central to our method. The entire process ensures a consistent and high-quality set of structured hints, and the scripts will be included in our code release.
B.3 Evaluation Benchmarks
@@
B.4 Motivation for Dataset Selection
Our choice of the DeepScaleR-Preview-Dataset for training is deliberate. This decision is motivated by two key factors. First, the dataset’s broad scope, encompassing problems of varying difficulty, provides the rich and diverse material necessary for our filtering strategy to be effective. Second, its successful application in prior research (Min et al., 2024) establishes it as a robust and relevant foundation for training advanced reasoning models.
For evaluation, our selection of benchmarks is designed for a rigorous and multifaceted assessment. The suite primarily consists of challenging competition-level benchmarks (AIME24/25, AMC, MATH, OlympiadBench, Minerva) and a standardized national exam (GaoKao2023), covering a broad spectrum of mathematical reasoning. Crucially, to measure out-of-distribution (OOD) generalization, we include the GPQA-Diamond benchmark (Rein et al., 2024). As GPQA consists of graduate-level questions whose style and domain are distinct from our training data, strong performance on this benchmark indicates that Scaf-GRPO fosters genuine reasoning skills rather than mere pattern memorization of the training distribution.
@@ B.4 Motivation for Dataset Selection
AIME24 American Invitational Mathematics Examination 2024 (AIME, 2024) 30
AIME25 American Invitational Mathematics Examination 2025 (AIME, 2025) 30
AMC American Mathematics Competitions 2023 (AMC, 2023) 25
MATH-500 A subset from the MATH test set (Hendrycks et al., 2021) 500
GaoKao2023en Chinese National College Entrance Exam 2023 (Chinese GaoKao Community, 2024) 385
OlympiadBench Math Olympiad-level problems (He et al., 2024) 675
Minerva A specialized dataset to evaluate quantitative and scientific reasoning abilities (Lewkowycz et al., 2022) 272
GPQA-Diamond Expert-level questions across biology, physics, and chemistry (Rein et al., 2024) 198
@@
Appendix C Experimental Setup Details
C.1 Computing Infrastructure
  • Framework: Our implementation is built upon the verl (0.4.1) framework (Sheng et al., 2025), a robust and efficient library designed for large-scale reinforcement learning with LLMs.
C.2 Hyperparameter Details
C.3 Implementation of Baseline Methods
@@
To compare against the dominant prefix-continuation paradigm, we train LUFFY (Yan et al., 2025) using its official public implementation and its original, recommended hyperparameter settings. To ensure a fair comparison of methodological effectiveness, we train it on our high-quality filtered dataset. This directly contrasts their off-policy guidance with our in-prompt scaffolding on identical data.
@@
For other leading methods like Simple-RL (Zeng et al., 2025) and Oat-Zero (Liu et al., 2025a), we do not perform any implementation or retraining. Instead, we evaluate their officially released, publicly available model checkpoints directly. All reported results for these models are obtained by running them through our unified evaluation pipeline, ensuring a consistent and fair comparison against established state-of-the-art work.
C.4 Evaluation Metrics Details
@@
  • For all mathematical reasoning benchmarks, we employ the “symeval” library (Tong, 2024), specifically its EvaluatorMathBatch module, to determine correctness. This approach moves beyond simple string comparison by using a sophisticated pipeline that combines regular expressions for robust answer extraction with SymPy for symbolic mathematical evaluation. This allows for the accurate verification of complex answers, including matrices, intervals, and symbolic expressions.
  • For the out-of-distribution GPQA-Diamond benchmark, we utilize the EleutherAI’s lm-evaluation-harness (Gao et al., 2024b) to ensure a fair and standardized assessment. This widely adopted framework provides a consistent testing environment for generative models. We use its implementation of the gpqa-diamond task to compute the pass@1 score, thereby maintaining metric consistency while leveraging a community-standardized evaluation harness for OOD generalization.
Appendix D Methodology Details
@@
D.1 The Progressive Exploration and Replacement Algorithm
@@
The algorithm leverages the pre-generated, four-level progressive hint structure, detailed in Appendix B.2. It systematically searches through the hint categories in order of decreasing abstraction ($H_{\text{knowledge}}\rightarrow H_{\text{planning}}\rightarrow H_{\text{solution}}$) and, within each category, through the four levels of increasing detail.
@@
Appendix E Prompt Design in Scaf-GRPO
@@
E.1 Hint Generation Prompt
@@
Prompt for hint injection **[ROLE & GOAL]** {internallinenumbers*} You are an expert AI assistant specializing in problem-solving methodology and knowledge engineering. Your task is to analyze a given problem and its ground-truth solution, and then generate a structured breakdown of the reasoning process with a high degree of granularity. **[INPUT]** I will provide you with a "Problem" and its "Ground-Truth Solution". **[INSTRUCTIONS]** {internallinenumbers*} Based on the provided input, you must generate exactly THREE components. For each component, you MUST generate **a minimum of 4 numbered items**. {internallinenumbers*} - If a natural breakdown results in fewer than 4 items, you must **subdivide the existing steps into more detailed, finer-grained sub-steps** to meet the requirement. For example, a single calculation step can be broken down into ’substituting values’, ’performing the operation’, and ’stating the result’. {internallinenumbers*} 1. **Planning Skeleton**: Extract a high-level planning skeleton. This should be a concise, ordered list of the key reasoning steps and the overall strategy used to reach the solution. Do not include detailed calculations, just the logical flow. Break it down into at least 4 detailed steps. {internallinenumbers*} 2. **Knowledge Components**: Identify at least 4 essential knowledge components (like facts, definitions, theorems, lemmas, or formulas) required to solve the problem. List each component clearly in a numbered list. {internallinenumbers*} 3. **Solution Breakdown**: Divide the original Ground-Truth Solution into a numbered list of semantically coherent steps or chunks.There should be at least 4 steps or chunks. Each item in the list should be a direct quote or a faithful summary of a part of the original solution text. **[OUTPUT FORMAT]** {internallinenumbers*} You MUST provide your response in the following structured format. Ensure each section contains at least 4 items. <PLANNING_SKELETON> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </PLANNING_SKELETON> <KNOWLEDGE_COMPONENTS> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </KNOWLEDGE_COMPONENTS> <SOLUTION_BREAKDOWN> 1. [Item 1] ... 4. [Item 4] ... (and more if applicable) </SOLUTION_BREAKDOWN> **[EXAMPLE]** --- BEGIN EXAMPLE --- # ... (Example Problem and Solution are the same) --- END EXAMPLE --- **[EXPECTED OUTPUT FOR THE EXAMPLE]** {internallinenumbers*} (This example now demonstrates the required granularity with >= 4 items per section) <PLANNING_SKELETON> {internallinenumbers*} 1. Identify the geometric shape (right-angled triangle) and the goal (find the hypotenuse). {internallinenumbers*} 2. Recall the relevant mathematical theorem that connects the sides of a right-angled triangle. {internallinenumbers*} 3. Formulate the equation by substituting the given side lengths (base and height) into the theorem. 4. Execute the arithmetic calculation to find the square of the hypotenuse. {internallinenumbers*} 5. Perform the final step of taking the square root to isolate the length of the hypotenuse. </PLANNING_SKELETON> <KNOWLEDGE_COMPONENTS> 1. Theorem: Pythagorean Theorem (a² + b² = c² for a right-angled triangle). {internallinenumbers*} 2. Definition: Hypotenuse (The longest side of a right-angled triangle, opposite the right angle). {internallinenumbers*} 3. Concept: Right-angled Triangle (A triangle with one angle measuring 90 degrees). {internallinenumbers*} 4. Mathematical Operation: Square Root (The inverse operation of squaring a number). </KNOWLEDGE_COMPONENTS> <SOLUTION_BREAKDOWN> {internallinenumbers*} 1. To find the hypotenuse of a right-angled triangle, we can use the Pythagorean theorem, which states that a² + b² = c², where a and b are the lengths of the two shorter sides (legs) and c is the length of the hypotenuse. 2. The given values are a = 4 cm and b = 3 cm. 3. Substituting these into the formula gives: 4² + 3² = 16 + 9 = 25. {internallinenumbers*} 4. This means c² = 25. Taking the square root of both sides results in c = 5 cm. </SOLUTION_BREAKDOWN> --- **[TASK]** {internallinenumbers*} Now, please process the following problem and solution, strictly following all instructions. **Problem**: {question} **Ground-Truth Solution**: {solution}
E.2 Inject Hints Prompt